12 of every 18 students ride the bus 9 of 15 live in hunt hill and 160 ride the bus. how many do not live in hunt hill
Answer by Theo(13342)

12 of every 18 students ride the bus.
160 ride the bus.

let x = number of students total.
12/18 * x = 160
solve for x to get x = 240.

there are 240 students in the school.

9/15 * 240 live in hunt hill.
9/15 * 240 = 144.

there are 144 students who live in hunts hill.

that means that 240 - 144 = 96 don't live in hunts hill.